# Evaluating LLMs on Sequential API Call Through Automated Test Generation

### v1

@@ III-B Trace Generation
III-C Program Generation
While the trace generated by TraceGenerator is already an executable program, we posit that incorporating control flow structures can further reflect the complexity in practical applications and allow for more comprehensive testing of the target LLMs. The injection of control flows also distinguishes our approach from prior work that relies solely on sequential execution [49, 31, 10, 24, 50].
The stateful design schema provided by TraceGenerator, makes it straightforward to add branches and other control flow constructs. One possible strategy involves first selecting a target point from all possible API call sites. When trace generation reaches this point, we create a copy of the TraceGenerator as the “else” TraceGenerator. Then we continue the “if” trace generation until it naturally terminates. The related data structures from the “if” run, such as the frequency recorder, will be passed to “else” TraceGenerator and complete its generation. This process is analogous to spawning a new thread, with both traces proceeding concurrently from the split. By repeating this procedure, we can build nested or multiple splits, but in our implementation, we limit a single split to keep the process computationally efficient and easy to manage.
@@ III-C Program Generation
III-D Instruction Translation
@@
III-E Scenario Selection and Oracle Generation
To demonstrate the effectiveness of the proposed framework, we implemented it in three representative usage scenarios.
  • Session Service in the EvoMaster Benchmark (EMB): EMB [25] is a well-established benchmark that provides a collection of web applications for researching fuzzing techniques. We focused on the session service, which manages cBioPortal user sessions through MongoDB. This application facilitates a full range of CRUD operations defined by OpenAPI specifications [45] and contains documentation totalling under 600 lines, making it fit for inclusion in LLM prompts. From this service, we selected five APIs.
  • PyTorch-Based Tensor Operations: PyTorch [26] is a widely used library for deep neural network computation. Its computation graphs require careful reasoning and precise management of tensor shapes across function calls. We chose six APIs for this scenario: four related to tensor manipulation and two focused on tensor computation (Linear and Conv2d). Related documentation was sourced directly from the official website.
  • ElevenLabs Model Context Protocol (MCP): MCP [51] is a newly introduced open protocol by Anthropic that aims to standardize interactions between LLMs and external tools. ElevenLabs [27] offers a set of AI tools for speech processing, such as text-to-speech, transcription, and outbound phone calls. From this collection, we chose six APIs and utilized the open-source documentation available within the protocol.
These three cases cover conventional RESTful API services, deep learning programming, and the latest LLM-driven tool usage. Together, they offer a broad and diverse set of playgrounds to evaluate the ability of LLMs in handling API calls and assess the effectiveness of our framework StateGen.
@@ III-E Scenario Selection and Oracle Generation
  • Tensor Operation: High-dimensional tensors are initialized at the start to ensure that results from computations (like Linear and Conv2d operations) and tensor manipulations are deterministic. At the evaluation stage, we verify both the shapes of tensors and the computation results.
@@
IV Evaluation
IV-A Experimental Setup
Benchmark Construction. By leveraging StateGen, a new benchmark, StateEval, is constructed to facilitate the evaluation of LLMs on sequential API calls. We first generate 60 programs for each of the three scenarios described in Section III-E. For each program, we set the number of APIs in the trace to five. [^2] When an if-else branch is present, the total number of API calls in the program (including both if and else branches) may exceed five. However, each individual execution trace, whether following the if or else branch, still contains five API calls. This number corresponds to the total number of APIs in the Session Service and, based on our observations, represents a moderate level of complexity that is neither trivial nor overly challenging for LLMs. For the insertion of control flow, we randomly select a point in each program to introduce an if-else branch; a do not insert option is also provided to accommodate programs with a single, linear trace. This approach is designed to maximize the diversity of the generated programs.
This study mainly revolves around four research questions. RQ1 investigates the effectiveness of our generation framework, StateGen. RQ2 compares our benchmark, StateEval, with existing executable benchmarks for code generation. RQ3 assesses the performance of both open-source and closed-source LLMs on our benchmark StateEval. RQ4 offers a detailed analysis of the errors made by these LLMs.
IV-B RQ1: Effectiveness of the Generation Framework StateGen
@@
For the first approach, we evaluate StateGen without the pair transition coverage guidance (StateGen-w/o-COV), as defined in Section III-B. This approach can also serve as an ablation study to assess the effectiveness of pair transition coverage guidance within our framework. In terms of the second, we implement an LLM-only approach, where LLMs are prompted to generate programs using the documentation for each application. In this setting, we retain the same constraints as in StateGen, such as restrictions on control flow and the number of API calls. We iteratively query GPT-4.1 until we obtain 40 programs for each application, ensuring the same number of test cases in StateEval. This LLM-only approach is inspired by recent studies highlighting the strong potential of LLMs for automated fuzzing [52, 53]. This approach is expected to reveal, without any auxiliary designs, the extent to which individual LLMs can generate diverse test cases for coding tasks with tool usage.
@@
where $M$ is the total number of unique API calls for a given application. It is important to note that the adjacent transition differs from the pair transition defined in Section III-B. While pair transition considers APIs bound by states, the adjacent transition only accounts for API calls that occur consecutively (that may not have data dependency). A higher ATC indicates that the generation method produces a broader range of local sequential structures, thereby implying a higher diversity in the generated test cases. This metric is particularly suitable for an initial comparative assessment of program structures generated by different approaches, including those that may not be fully executable or easily analyzable with pairwise transitions.
@@
We present the results in Figure 3. Our results show that StateGen achieves higher coverage at a faster rate compared to the other two baselines. In particular, StateGen converges more quickly than the random baseline and is able to discover new adjacent API calls, such as those found in the Session Service. In contrast, the LLM-only approach demonstrates a significantly lower coverage rate than both StateGen and the random baseline, which indicates a lower diversity in the generated test cases. This finding aligns with recent theoretical work suggesting that LLM-based sampling can be inherently biased [54], making it less suitable to serve as an end-to-end fuzzing engine.
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.5pt, leftmargin=1pt, rightmargin=1pt, innerleftmargin=5pt, innerrightmargin=5pt, topline=true, bottomline=true, skipabove=10pt ] Answer to RQ1: StateGen demonstrates superior effectiveness in generating sequential API calls, achieving higher coverage and faster convergence compared to other baselines.
IV-C RQ2: Comparison of StateEval with Existing Benchmarks
  • HumanEval [19, 5] is one of the most widely adopted benchmarks, comprising 164 programming problems designed to evaluate LLMs’ capabilities in language comprehension and program synthesis.
  • DS-1000 [21] includes 1000 data science problems collected from StackOverflow spanning seven Python libraries.
  • BigCodeBench [20] is a recently introduced benchmark that evaluates LLMs’ capability of handling diverse function calls and understanding complex instructions.
  • Berkeley Function Calling Leaderboard (BFCL) [10] is a popular leaderboard with APIBench, aiming to perform a comprehensive evaluation of the LLM’s proficiency in invoking functions and interacting with tools. We select parallel and multiple executable subsets of BFCL, which have 140 samples.
@@ IV-C RQ2: Comparison of StateEval with Existing Benchmarks
Two metrics, Path Depth [55] and Binding Count [56, 57], are applied to further characterize the complexity of these benchmarks from both data and structural dependency perspectives. Path Depth quantifies the maximum length of a data-dependent chain among API calls, where the output of one call serves as the input to another. A higher Path Depth indicates more deeply nested data dependencies and longer sequential reasoning chains. Additionally, Binding Count captures the number of shared variables that create data bindings across API calls (i.e., where one call’s output is reused as another’s input). A higher Binding Count suggests a greater density of data interconnections between API calls within a program. Together, these two metrics reflect the degree of inter-dependency among function calls within a program. To compute both metrics, we perform data flow analysis and construct call graphs to identify relevant dependencies. Table I shows that StateEval exhibits substantially higher values for both Path Depth and Binding Count, indicating that it does not simply stack multiple unrelated API calls. Instead, it constructs meaningful dependencies across calls, resulting in more interdependent program structures. It is also worth noting that BFCL yields zero for both metrics, as it typically prompts LLMs to generate API calls that do not share any parameters.
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.5pt, leftmargin=1pt, rightmargin=1pt, innerleftmargin=5pt, innerrightmargin=5pt, topline=true, bottomline=true, skipabove=10pt ] Answer to RQ2: High-level statistics suggest that StateEval is well-suited for evaluating LLMs’ capacity to understand complex instructions and produce multi-API calls with rich interdependencies.
@@
Qwen2.5-Coder [58] 32B Yes
Llama-4-Scout (MoE) [59] 17B (/109B) Yes
Gemini-2.5-Flash [60] Unknown No
GPT-4.1-nano [61] Unknown No
GPT-4.1 [62] Unknown No
TABLE II: Open-source and closed-source LLMs used in our evaluation.
IV-D RQ3: Evaluation of LLM Performance
@@
We evaluate in total five LLMs, including two open-source models and three commercial options, using our StateEval. The candidate LLMs are shown in Table II. Among these models, Qwen2.5-Coder from the Qwen Team and Llama-4-Scout from Meta are high-performance open-source LLMs that are publicly available and have been widely studied in the literature. For both models, we use the instruct versions. The remaining models, Gemini from Google and the GPT-4.1 family from OpenAI, are commercial LLMs released after April 2025, representing the most recent advancements.
We use pass@1 rates to assess the performance of target models. As shown in Figure 4, closed-source models exhibit better capabilities in addressing API-oriented tasks. Specifically, GPT-4.1 achieves the highest pass@1 rate at 56%, significantly outperforming the open-source models (25% for Qwen2.5-Coder and 34% for Llama-4-Scout). Gemini-2.5-Flash ranks second among all models with a pass@1 rate of 45%. In contrast, the cost-efficient GPT-4.1-nano achieves a pass@1 rate of 35%, which is slightly lower than that of open-source Llama-4-Scout.
@@
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.5pt, leftmargin=1pt, rightmargin=1pt, innerleftmargin=5pt, innerrightmargin=5pt, topline=true, bottomline=true, skipabove=10pt ] Answer to RQ3: There remains substantial room for improvement for LLMs on StateEval, particularly among open-source models. Although all the generated programs are unlikely to appear in LLMs’ training data, greater familiarity with relevant APIs can still significantly improve their performance.
IV-E RQ4: Error Analysis of LLMs
@@
We conduct a detailed analysis of the errors made by LLMs on StateEval, categorizing them into three types: (1) Syntax errors, where the generated code contains invalid syntax and cannot be parsed; (2) Execution errors, which occur when a program starts running but terminates unexpectedly, often due to issues with data handling, control flow construction, API selection, or API parameter reasoning; and (3) Result errors, where the program executes without interruption but produces an incorrect final result or side effect, such as failing to update a remote database. The distribution of these error types is shown in Figure 5.
@@
In general, execution and result errors are the most prevalent. Execution errors are especially common in Session Service and Tensor Operation tasks. In Session Service, LLMs may attempt to access non-existent data items, resulting in None values and potentially causing downstream errors such as IndexError when operating on empty lists. In Tensor Operation, improper handling of API calls can result in incompatible tensor shapes, causing downstream failures (e.g., RuntimeError encountered for incompatible matrix multiplication in the Linear layer). Conversely, most errors in Elevenlabs MCP are result errors. This is because tool calling in this application is less restrictive regarding input types, but incorrect API usage can still lead to state transition errors, which are only detected during final result checking. For example, the LLMs may fail to place phone calls using the designated voice or may select incorrect parameters during text-to-speech conversion, causing the final speech to not be as expected.
Although relatively rare, some syntax errors still occur, primarily with Qwen2.5-Coder. Many of these involve string manipulation, particularly f-strings, and inconsistencies in the output code format compared to what is specified in the instructions.
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.5pt, leftmargin=1pt, rightmargin=1pt, innerleftmargin=5pt, innerrightmargin=5pt, topline=true, bottomline=true, skipabove=10pt ] Answer to RQ4: The inclusion of a variety of application types allows StateEval to assess how well LLMs interpret multi-step instructions, understand API documentation, and manage complex interdependencies between API calls. As a result, StateEval can provide valuable insights into both the strengths and limitations of LLMs in handling different types of application logic.

### v2

@@ III-C Trace Generation
For a more concrete illustration, suppose there are two valid API candidates following Snippet 1, linear and conv2d, both of which accept r_1 as input. Using r_1 as the linking variable, StateGen identifies two possible transitions: (cat, linear) and (cat, conv2d). StateGen then consults the frequency recorder and observes that (cat, conv2d) has been selected less often than (cat, linear), and therefore assigns a higher selection probability to the former. If conv2d is chosen and appended to the trace (as shown in Snippet 2), all relevant state information is updated, and the newly generated variable r_2 is recorded by StateGen.
Code Snippet 2: Second API
    1  ut1 = torch.load("user1")  2  ut2 = torch.load("user2")  3 # API 1: concatenate the tensors from 4 # initilized variables ut1 and ut2  5 # along dimension 0 6  r_1 = torch.cat((ut1, ut2), 0)  7 # API 2: A legal call for conv2d based on 8 # r_1 and ut1 9  r_2 = F.conv2d(r_1, ut1, stride=2,  10   padding=0, dilation=1)
III-D Program Generation
While the trace generated by TraceGenerator is already an executable program, we posit that incorporating control flow structures can further reflect the complexity in practical applications and allow for more comprehensive testing of the target LLMs. The injection of control flows also distinguishes our approach from prior work that relies solely on sequential execution [li-etal-2023-api, basu2024nestful, patil2024gorilla, wu2024comprehensive, sheng2024measuring].
The stateful design schema provided by TraceGenerator, makes it straightforward to add branches and other control flow constructs. One possible strategy involves first selecting a target point from all possible API call sites. When trace generation reaches this point, we create a copy of the TraceGenerator as the “else” TraceGenerator with all local variables. Then we continue the “if” trace generation until it naturally terminates. The related global data structures from the “if” run, such as the frequency recorder, will be passed to “else” TraceGenerator and complete its generation. This process is analogous to spawning a new thread, with both traces proceeding concurrently from the split. By repeating this procedure, we can build nested or multiple splits, but in our implementation, we limit a single split to keep the process computationally efficient and easy to manage.
@@ III-D Program Generation
We illustrate this process using the example in Snippet 3. Suppose the condition insertion point is placed immediately after the second API call in Snippet 2. In this case, StateGen selects r_2 as the left-hand-side variable of the if statement, since it is the most recently produced variable, and inserts a placeholder (condition_1) for the condition, which will be initialized at the start of the program. After this point, two TraceGenerator proceed in parallel, each maintaining its own data structures along its respective branch until the final API is generated. The sink variable RESULT then aggregates the outputs from both branches, which are compared against ground truth during evaluation.
Code Snippet 3: Full Example
    1 # Initialization Area 2  ut1 = torch.load("user1")  3  ut2 = torch.load("user2")  4 ## placeholder for the if-condition 5  condition_1 = 128  6 # API 1 7  r_1 = torch.cat((ut1, ut2), 0)  8 # API 2 9  r_2 = F.conv2d(r_1, ut1, stride=2,  10   padding=0, dilation=1)  11 # --->If Condition Insertion 12  if r_2.shape[1] == condition_1:  13  # Traces Generated by TraceGenerator 1 14  # ... 15   RESULT = ...  16  else:  17  # Traces Generated by TraceGenerator 2 18  # ... 19   RESULT = ...
III-E Instruction Translation
@@
III-F Oracle Generation
@@
  • Tensor Operation: High-dimensional tensors are initialized at the start (i.e., loading from disk instead of randomly created) to ensure that results from computations (such as Linear and Conv2d operations) and tensor manipulations are deterministic. At the evaluation stage, we verify both the shapes of tensors and the computation results.
@@
IV Evaluation
IV-A Experimental Setup
Benchmark Construction. By leveraging StateGen, a new benchmark, StateEval, is constructed to facilitate the evaluation of LLMs on sequential API calls. We first generate 60 programs for each of the three scenarios described in Section III-F. For each program, we set the number of APIs in the trace to five. [^3] When an if-else branch is present, the total number of API calls in the program (including both if and else branches) may exceed five. However, each individual execution trace, whether following the if or else branch, still contains five API calls. This number corresponds to the total number of APIs in the Session Service and, based on our observations, represents a moderate level of complexity that is neither trivial nor overly challenging for LLMs. Using too many APIs risks out-of-memory errors during evaluation of locally deployed LLMs, while using too few limits the program’s potential complexity. For the insertion of control flow, we randomly select a point in each program to introduce an if-else branch; a do not insert option is also provided to accommodate programs with a single, linear trace. This approach is designed to maximize the diversity of the generated programs.
This study mainly revolves around four research questions. RQ1 investigates the effectiveness of our generation framework, StateGen. RQ2 compares our benchmark, StateEval, with existing executable benchmarks for code generation. RQ3 assesses the performance of both open-source and closed-source LLMs on our benchmark StateEval. RQ4 presents a quantitative analysis of the symptoms of LLM errors, and RQ5 follows with case studies that illustrate representative root causes behind these errors.
IV-B RQ1: How effective is the generation framework StateGen?
@@
For the first approach, we evaluate StateGen without the pair transition coverage guidance (StateGen-w/o-COV), as defined in Section III-C. This approach can also serve as an ablation study to assess the effectiveness of pair transition coverage guidance within our framework. In terms of the second, we implement an LLM-only approach, where LLMs are prompted to generate programs using the documentation for each application. In this setting, we retain the same constraints as in StateGen, such as restrictions on control flow and the number of API calls. We iteratively query GPT-4.1 until we obtain 40 programs for each application, ensuring the same number of test cases in StateEval. This LLM-only approach is inspired by recent studies highlighting the strong potential of LLMs for automated fuzzing [xia2024fuzz4all, lemieux2023codamosa]. This approach is expected to reveal, without any auxiliary designs, the extent to which individual LLMs can generate diverse test cases for coding tasks with tool usage.
@@
where $M$ is the total number of unique API calls for a given application. It is important to note that the adjacent transition differs from the pair transition defined in Section III-C. While pair transition considers APIs bound by states, the adjacent transition only accounts for API calls that occur consecutively (that may not have data dependency). A higher ATC indicates that the generation method produces a broader range of local sequential structures, thereby implying a higher diversity in the generated test cases. This metric is particularly suitable for an initial comparative assessment of program structures generated by different approaches, including those that may not be fully executable or easily analyzable with pairwise transitions.
We present the results in Figure 3. Our results show that StateGen achieves higher coverage at a faster rate compared to the other two baselines. In particular, StateGen converges more quickly than the random baseline and is able to discover new adjacent API calls, such as those found in the Session Service. In contrast, the LLM-only approach demonstrates a significantly lower coverage rate than both StateGen and the random baseline, which indicates a lower diversity in the generated test cases. This finding aligns with recent theoretical work suggesting that LLM-based sampling can be inherently biased [sivaprasad2025theoryllmsamplingdescriptive], making it less suitable to serve as an end-to-end fuzzing engine.
Answer to RQ1: StateGen demonstrates superior effectiveness in generating sequential API calls, achieving higher coverage and faster convergence compared to other baselines.
IV-C RQ2: How does StateEval compare with existing benchmarks?
  • HumanEval [chen2021codex, liu2023your] is one of the most widely adopted benchmarks, comprising 164 programming problems designed to evaluate LLMs’ capabilities in language comprehension and program synthesis.
  • DS-1000 [lai2023ds] includes 1000 data science problems collected from StackOverflow spanning seven Python libraries.
  • BigCodeBench [zhuo2025bigcodebench] is a recently introduced benchmark that evaluates LLMs’ capability of handling diverse function calls and understanding complex instructions.
  • Berkeley Function Calling Leaderboard (BFCL) [patil2024gorilla] is a popular leaderboard with APIBench, aiming to perform a comprehensive evaluation of the LLM’s proficiency in invoking functions and interacting with tools. We select parallel and multiple executable subsets of BFCL, which have 140 samples.
@@ IV-C RQ2: How does StateEval compare with existing benchmarks?
Two metrics, Path Depth [zhang2011flow] and Binding Count [beyer2010simple, peitek2021program], are applied to further characterize the complexity of these benchmarks from both data and structural dependency perspectives. Path Depth quantifies the maximum length of a data-dependent chain among API calls, where the output of one call serves as the input to another. A higher Path Depth indicates more deeply nested data dependencies and longer sequential reasoning chains. Additionally, Binding Count captures the number of shared variables that create data bindings across API calls (i.e., where one call’s output is reused as another’s input). A higher Binding Count suggests a greater density of data interconnections between API calls within a program. Together, these two metrics reflect the degree of inter-dependency among function calls within a program. To compute both metrics, we perform data flow analysis and construct call graphs to identify relevant dependencies. Table I shows that StateEval exhibits substantially higher values for both Path Depth and Binding Count, indicating that it does not simply stack multiple unrelated API calls. Instead, it constructs meaningful dependencies across calls, resulting in more interdependent program structures. It is also worth noting that BFCL yields zero for both metrics, as it typically prompts LLMs to generate API calls that do not share any parameters.
Answer to RQ2: High-level statistics suggest that StateEval is well-suited for evaluating LLMs’ capacity to understand complex instructions and produce multi-API calls with rich interdependencies.
TABLE II: Open-source and closed-source LLMs used in our evaluation.
@@
Qwen2.5-Coder [hui2024qwen2] 32B Yes
Llama-4-Scout (MoE) [meta2025llama4] 17B (/109B) Yes
Gemini-2.5-Flash [deepmind2025gemini25flash] Unknown No
GPT-4.1-nano [openai2025gpt41nano] Unknown No
GPT-4.1 [openai2025gpt41] Unknown No
IV-D RQ3: How well do existing LLMs perform?
@@
We evaluate in total five LLMs, including two open-source models and three commercial options, using our StateEval. The candidate LLMs are shown in Table II. Among these models, Qwen2.5-Coder from the Qwen Team and Llama-4-Scout from Meta are high-performance open-source LLMs that are publicly available and have been widely studied in the literature. For both models, we use the instruction-tuned versions. The remaining models, Gemini from Google and the GPT-4.1 family from OpenAI, are commercial LLMs released after April 2025, representing the most recent advancements.
We use pass@1444Although we use the term pass@1, our evaluation includes checks beyond standard assertion-based testing used in other code-generation benchmarks. In addition to verifying functional outputs, we also validate underlying states—such as server-side transitions in MCP workflows or data updates in remote databases. For simplicity, we use pass@1 to indicate that a generated code snippet satisfies all checks defined in our benchmark. rates to assess the performance of target models. As shown in Figure 4, closed-source models exhibit better capabilities in addressing API-oriented tasks. Specifically, GPT-4.1 achieves the highest pass@1 rate at 56%, significantly outperforming the open-source models (25% for Qwen2.5-Coder and 34% for Llama-4-Scout). Gemini-2.5-Flash ranks second among all models with a pass@1 rate of 45%. In contrast, the cost-efficient GPT-4.1-nano achieves a pass@1 rate of 35%, which is slightly lower than that of open-source Llama-4-Scout.
@@
Answer to RQ3: There remains substantial room for improvement for LLMs on StateEval, particularly among open-source models. Although all the generated programs are unlikely to appear in LLMs’ training data, greater familiarity with relevant APIs can still significantly improve their performance.
IV-E RQ4: What are the symptoms of LLM errors?
We first conducted a quantitative analysis of the errors made by LLMs on StateEval, categorizing them into three types:
@@ IV-E RQ4: What are the symptoms of LLM errors?
(1) Syntax errors, where the generated code contains invalid syntax and cannot be parsed;
(2) Execution errors, which occur when a program starts running but terminates unexpectedly, often due to issues with data handling, control flow construction, API selection, or API parameter reasoning;
(3) Result errors, where the program executes without interruption but produces an incorrect final result or side effect, such as failing to update a remote database. The distribution of these error types is shown in Figure 5.
@@
In general, execution and result errors are the most prevalent. Execution errors are especially common in Session Service and Tensor Operation tasks. In Session Service, LLMs may attempt to access non-existent data items, resulting in None values and potentially causing downstream errors such as IndexError when operating on empty lists. In Tensor Operation, improper handling of API calls can result in incompatible tensor shapes, causing downstream failures (e.g., RuntimeError encountered for incompatible matrix multiplication in the Linear layer). Conversely, most errors in Elevenlabs MCP are result errors. This is because tool calling in this application is less restrictive regarding input types, but incorrect API usage can still lead to state transition errors, which are only detected during final result checking. For example, the LLMs may fail to place phone calls using the designated voice or may select incorrect parameters during text-to-speech conversion, causing the final speech to not be as expected. On the other hand, syntax errors are relatively uncommon, with Qwen2.5-coder exhibiting them most frequently. This is expected, as Qwen2.5-coder is the smallest model among the weaker open-source ones and can occasionally struggle to adhere to prompt instructions or produce syntax-correct code.
In summary, the test oracles in StateEval, together with the carefully selected scenarios, enable a relatively holistic evaluation of an LLM’s ability to generate code involving tool invocations across a broad range of dimensions.
Answer to RQ4: On StateEval, failures cluster into three types: syntax (unparsable code), execution (crashes from data/control/API or parameter mistakes), and result (runs but yields a wrong outcome). Among these errors, execution and result are dominant while syntax is rare. By spanning diverse, stateful tasks with strict oracles, StateEval can reliably evaluate LLMs across different scenarios.
IV-F RQ5: What are the representative root causes of LLM errors?
[…]
Example 1: Qwen2.5-coder error on Session Service. Here we show generated code (left) compared with the correct reference code (right).
# Generated Code def create_session(source, session_type, data): url = f"{BASE_URL}/{source}/{session_type}" response = requests.post(url, json=data) response.raise_for_status() return response.json() # ... # Correct Code def create_session(source, session_type, data): url = f"{BASE_URL}/api/sessions/{source}/{session_type}" response = requests.post(url, json=data) response.raise_for_status() return response.json() # ...
[…]
Example 2: GPT-4.1-nano error on Elevenlabs MCP. Here we show generated code (left) compared with the correct reference code (right).
# Generated Code search_results = search_voice_library( search_category=user_variable_0.get(’search_category’), search_gender=user_variable_0.get(’search_gender’) ) if search_results: first_voice = search_results[0] voice_name = first_voice.get(’voice_name’) isolated_speech = isolate_audio(input_speech=user_variable_0_load) # ... Omit for simplicity # Instruction: Transform the isolated audio (from # user_variable_0_load) into speech transformed_speech = text_to_speech( text=user_variable_0_load, voice_name=voice_name ) # ... # Correct Code search_results = search_voice_library( search_category=user_variable_0.get(’search_category’), search_gender=user_variable_0.get(’search_gender’) ) if search_results: first_voice = search_results[0] voice_name = first_voice.get(’voice_name’) isolated_speech = isolate_audio(input_speech=user_variable_0_load) # ... Omit for simplicity # Instruction: Transform the isolated audio (from # user_variable_0_load) into speech transformed_speech = speech_to_speech( input_speech=isolated_speech, voice_name=voice_name ) # ...
Finally, Example 3 illustrates the third type of failure (i.e., mismanaging implicit states across consecutive API calls) in the context of tensor operations on Llama4-scout. The model was instructed to perform a sequence of tensor manipulations. However, in the generated code, the model incorrectly inferred the shape of the conv_result_2 matrix and inappropriately invoked view for reshaping. Consequently, this caused a shape mismatch error (”mat1 and mat2 shapes cannot be multiplied”) in the linear computation. This mistake caused the code to fail at runtime. Addressing this challenge may require strengthening the model’s mathematical reasoning capabilities.
Example 3: Llama4-scout error on Tensor Operation. Here we show generated code (left) compared with the correct reference code (right).
# Generated Code # ... # Output shape: torch.Size([5, 42, 42, 21]) transposed_tensor = torch.transpose(user_tensor_0, 1, 3) if list(user_tensor_0.shape) == user_constant_0: # ... Omit for simplicity. else: # ... Omit for simplicity. # Output shape: torch.Size([5, 24, 42, 21]) conv_result_2 = F.conv2d(transposed_tensor, user_weight_2, stride=1, padding=’same’, dilation=1) # user_weight_3 shape: torch.Size([21, 12]) linear_result = F.linear(conv_result_2.view(-1, 42), user_weight_3) # ... # Correct Code # ... # Output shape: torch.Size([5, 42, 42, 21]) transposed_tensor = torch.transpose(user_tensor_0, 1, 3) if list(user_tensor_0.shape) == user_constant_0: # ... Omit for simplicity. else: # ... Omit for simplicity. # Output shape: torch.Size([5, 24, 42, 21]) conv_result_2 = F.conv2d(transposed_tensor, user_weight_2, stride=1, padding=’same’, dilation=1) # user_weight_3 shape: torch.Size([21, 12]) linear_result = F.linear(conv_result_2, user_weight_3) # ...
Answer to RQ5: Our case studies reveal three representative root causes behind LLM failures in generating code with tool invocations. First, models can sometimes misinterpret API documentation, overlooking required conditions. Second, in certain scenarios, they fail to follow users’ instructions, leading to incorrect operations. Third, they may struggle with interdependent API calls, producing inconsistent data flows or broken dependencies.